OpenStudy (anonymous):

Given the exponential function A(x) = P(1 + r)x, what value for r will make the function a growth function? @shamil98 @ranga

OpenStudy (shamil98):

Don't know this one.

OpenStudy (anonymous):

@idku

OpenStudy (agent0smith):

If the 1+x is greater than 1, then it's a growth function. For growth, 1+x > 1 If 1+x is less than 1, it's a decay function. For decay, 1+x < 1

OpenStudy (anonymous):

r = −2.1 r = −0.1 r = 0 r = 2.1 these are the options @agent0smith

OpenStudy (shamil98):

Which of those values make 1 + x > 1 true?

OpenStudy (agent0smith):

Well read what I wrote above, and see if you can figure it out!

OpenStudy (anonymous):

is it 0 @agent0smith

OpenStudy (agent0smith):

Just realized it should be 1+r not 1+x, but still. Shouldn't be hard to decide which numbers will make it more than 1... you can do a few examples to help if you don't understand what we wrote above: if r = 0.5, then 1 + 0.5 = 1.5 if r = -0.3 1 - 0.3 = 0.7

OpenStudy (agent0smith):

if r=0 1 + 0 = 1

OpenStudy (anonymous):

-2.1

OpenStudy (agent0smith):

Are you just guessing?

OpenStudy (agent0smith):

You need 1 + r to be bigger than 1, right? If r=-2.1, then what is 1 - 2.1 equal to?

OpenStudy (anonymous):

ohh so it would be 2.1

OpenStudy (ranga):

Yes, r should be positive to make (1+r) > 1 and that will be a growth function.